OpenStudy (anonymous):

I'm just wondering because I'm just starting to learn about logarithms. Why & how are they important?

OpenStudy (campbell_st):

the only real place logarithms are used now are in engineering. before calculators and computers there were used to make complex arithmetic easier.

OpenStudy (anonymous):

They're useful for solving many types of equations involving exponents. For example: \[2^x = 10\] can be solved by taking the log (to any base) of both sides to get: \[x \log 2 = \log 10\] \[x = \log 10 / \log 2\] They were also used to perform multiplication and division before calculators by adding logs.

OpenStudy (campbell_st):

lol... when was the last time you were walking down the street and someone said Ok... if \[2^x = 13\] whats the value of x...

OpenStudy (anonymous):

His question wasn't "when in my everyday life will logarithms help me." He asked how they're useful. And the are very useful in mathematics and to people who care about mathematics.

OpenStudy (campbell_st):

@vinnv226 being able to relate mathematics to the real world is the key to any mathematics that you undertake...

OpenStudy (anonymous):

@campbell_st True. But not assuming that any mathematics you take is going to be useful someday because someone is going to come up to you on the street and ask. When has anyone come up to you on the street and asked you what 2+3 is? Does that make addition useless and unimportant?